### v4

Abstract
This paper compared physics-informed neural network (PINN), conventional neural network (NN) and traditional numerical discretization methods on solving differential equations (DEs) through literature investigation and experimental validation. We focused on the soft-constrained PINN approach and its mathematical framework and computational flow for solving Ordinary DEs and Partial DEs (ODEs/PDEs) was formalized. The working mechanism and its accuracy and efficiency were experimentally verified by solving typical linear and non-linear oscillator ODEs. We demonstrate that the DeepXDE-based implementation of PINN is not only light code and efficient in training, but also flexible across CPU/GPU platforms. PINN greatly reduces the need for labeled data: when the nonlinearity of the ODE is weak, a very small amount of supervised training data plus a few unsupervised collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (with initial values) are needed for first- or second-order ODEs, respectively. Strongly nonlinear ODE also require only an appropriate increase in the number of training points, which still has significant advantages over conventional NN. We also find that, with the aid of collocation points and the use of physical information, PINN has the ability to extrapolate data outside the time domain of the training set, and especially is robust to noisy data, thus with enhanced generalization capabilities. Training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. The soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law. Furthermore, PINN can also be used for stiff ODEs, PDEs, and other types of DEs, and is becoming a favorable catalyst for the era of Digital Twins.
keywords:
PINN; Oscillator ODEs; Soft-constrained; DeepXDE; Minimalistic Data; Noise; Non-linear; Regularization of Conservation of Energy
1 Introduction
1.1 Why PINN: In Small Data Situations where Neural Networks can be Improved
With the explosive growth of available data and computational resources, deep learning have achieved superior human performance in several tasks such as image recognition and chess playing [1, 2]. Its success relies first and foremost on big data and thus suffers from overfitting [3], poor generalization performance and interpretability issues [4]. On the other hand, in real complex scientific and engineering (e.g., physical, biochemical, medical) practices [5, 6, 7], data are often generated through expensive experiments or large-scale simulations, which are often sparse and noisy. Sometimes the cost of obtaining data is prohibitively high, and large amounts of experimental data are not even available at all [8, 9].
In these small data situations, most neural network methods (including convolutional and recurrent neural networks) lack robustness and do not provide convergence guarantees [10]. Thus, we inevitably face the challenge of predicting or making decisions with partial information. The good news is that there is a great deal of prior knowledge (i.e., the culmination of previous wisdom) in these system modeling cases that has not been fully exploited in traditional neural networks. Deep neural networks are highly expressive, thus neglecting to utilize a priori information or knowledge, yet also leading to an overemphasis and reliance on data as well as computational power.
Raissi et al. [10] proposed the idea of utilizing physical information or prior knowledge in the learning process of neural networks, i.e., Physics-informed Neural Network (PINN). This is a new class of general-purpose function approximators that inherits the strong expressive power of neural networks. It is able to encode any of the fundamental laws of physics that govern a given data set, which is often described in terms of differential equations. In other words, it is the process of improving the performance of learning algorithms by utilizing a priori knowledge derived from our observations, experiences, and physical or mathematical understanding of the world, which includes reducing data redundancy, improving generalization performance, increasing computational efficiency, and also increasing robustness and interpretability.
PINN incorporates physical information and knowledge into network topology or computational processes as model priors, in a structured, modular neural network learning architecture. There are two specific implementations of PINN: the approaches of hard constraints and soft constraints. Hard constraints [11] generally ensure that physical knowledge or laws (e.g., differential equations, symmetries, conservation laws [12, 13, 14]), as well as boundary and initial conditions [15], are strictly adhered to by hard-coding or embedding them directly as part of the neural network architecture or computational process, which needs to be approximated by a specific network design or training strategy [16]. The soft constraints discussed in this paper, on the other hand, are indirectly achieved by adding data and governing equation residual or regularization terms to the loss function, which is easier to implement. This approach can be considered as a specific use case for multi-task learning. Soft constraints allow some flexibility or tolerance in satisfying physical laws or conditions, but require balancing predictive performance and physical accuracy by carefully adjusting the various weighting factors in the loss function.
The key property of PINN is that it generalizes well even with only a small amount of supervised or labeled training data. Furthermore, these prior knowledge or constraints can yield more interpretable learning methods. These methods remain robust to data imperfections (e.g., noise, missing or outliers) and provide accurate and physically consistent predictions that can even be used for extrapolation tasks [8].
1.2 PINN as a Complement or Alternative to Numerical Discretization
Differential equations such as ordinary differential equations (ODEs) and partial differential equations (PDEs) are important tools for expressing the laws of nature in science and engineering. Real-world physical systems can be modeled using these differential equations based on different domain-specific assumptions and simplifications. These models can be used to approximate the behavior of these systems [17]. The common representative numerical discretization methods for solving ODEs/PDEs are the Runge-Kutta method [18] and the finite element method (FEM) [19], respectively. From a view of interpretability, these methods are well understood due to the simplicity of the underlying approximation, with readily available error estimates as well as convergence and stability guarantees.
Great progress has been made in solving differential equations (ODE/PDE) by numerical discretization in order to simulate various types of field problems, and its theoretical foundations are well established, with the advantages of high efficiency, high accuracy, and good stability. However, there are still some bottlenecks that severely limit its application [8]: (i) the curse of dimensionality problem, (ii) the mesh generation is still complicated, (iii) there are difficulties in merging experimental data, still unable to seamlessly incorporate noisy data into the existing algorithms, and (iv) it is not possible to solve the high-dimensional problems controlled by the parameterized PDEs, and so on. For example, an obvious disadvantage of FEM is that it relies on spatial discretization (spatial meshes plus large polynomial bases) and suffers from the curse of dimensionality: it is already difficult to use it in three dimensions, let alone for higher dimensional problems. Moreover, certain nonlinear and non-smooth PDEs are very difficult to discretize, e.g., due to general non-smooth behavior or singularities, which usually need to be solved on a very fine mesh. Since FEM is a mesh-based solver, solving over certain irregular domains requires customized methods that are challenging to design and solve [17]. In addition, solving inverse problems (e.g., for inferring material properties in functional materials or discovering missing physics in reaction transport) is often very expensive and requires complex formulas, new algorithms, and sophisticated computer codes. Moreover, solving realistic physical problems with missing, gapped, or noisy boundary conditions by conventional methods is currently impossible [8].
Deep learning methods (e.g., PINN [10], deep Ritz method [20], deep Galerkin method [21]) have achieved great success in solving high-dimensional problems and have recently become a common method for numerical solutions of various PDEs. They have the potential to overcome some of the challenges faced by the numerical discretization methods described above: (i) By utilizing automatic differentiation [22], the need for discretization is eliminated, giving the advantage of being mesh-free. (ii) Neural networks are able to represent more general functions than finite element bases, can break the curse of dimensionality [23], providing a promising direction for solving high-dimensional PDEs. (iii) Although training neural networks (the non-convex optimization problem) may become computationally intensive compared to numerical solvers such as FEM, it is very effective and efficient in evaluating new data samples. (iv) Its ability and flexibility to integrate data and knowledge provides a scale-able framework for dealing with problems with imperfect knowledge or limited data. However, training deep neural networks requires a lot of data, which is not always available for related scientific problems, as mentioned earlier.
Among them, PINN serves as a potentially disruptive technology that integrates physical information and knowledge into network topology or computational process. PINN can obey any symmetry, in-variance, or conservation principle derived, for example, from the laws of physics governing observational data, modeled by usual time-varying and nonlinear PDEs [10, 24]. This simple yet powerful construction allows us to address a wide range of problems in computer science, leading to the development of new data-efficient and physics-informed learning machines, new numerical solvers for ODEs/PDEs, and new data-driven methods for model inversion and system identification. It should be emphasized that PINN is not intended to replace FEM. PINN is particularly effective in solving, for example, the hypothetical and inverse problems described in this section; and for forward problems that do not require any data assimilation, existing grid-based numerical solvers currently outperform PINN [8, 17]. Table 1 summarizes the advantages and disadvantages of the three types of methods, i.e., traditional numerical discretization, conventional neural networks (NN), and PINN for solving differential equations.
1.3 Characteristics of the PINN Method Revisit
First of all, PINN has the unique advantage of achieving strong generalization in small data regime. By enforcing or embedding physics, the neural network model is effectively constrained to a lower dimensional manifold and therefore requires only a small amount of data for training. PINN is not only capable of interpolation but also extrapolation. A natural question here is how minimally dependent is PINN on data?
It is important to emphasize that, unlike any classical numerical method for solving differential equations (mesh-based or discrete step-size methods such as 4th-order Runge-Kutta, FDM, FEM), the predictions of PINN are obtained without any discretization of the spatio-temporal domain and it is able to break the curse of dimensionality [23, 25]. PINN does not need to deal with prohibitively small step sizes, so it can easily handle irregular and moving domain problems [17] and scales well in higher dimensions. Automatic differentiation is a significant advantage of PINN over other classical numerical methods for solving DEs, which utilizes the chain rule to back-propagate through the network and compute the derivatives, being able to accurately evaluate the differential operators at the configuration points with machine accuracy [26]. It does not depend on the discrete grid of the domain, but for which the sampling method becomes more important [27, 28].
PINN seamlessly integrates data with mathematical-physical models, even in partially understood, uncertain situations of imperfect data. Due to the inherent smoothness or regularity of the PINN formulation, it is possible to find meaningful solutions even if the problem assumptions or models are incomplete [9]. Moreover, PINN can directly handle nonlinear problems (an inherent property of neural networks that are good at handling nonlinear problems) without a priori assumptions, linearization, or local time stepping [10]. PINN is effective and efficient in dealing with the ill-posed and inverse problems, e.g., forward and inverse problems with no initial or boundary conditions specified, or problems where some parameters in the PDEs are unknown [8], when classical numerical methods may fail.
Therefore, based on the literature ivestigation, this work aims to further demonstrate the main characteristics and working mechanism of soft-constrained PINN by solving typical oscillator ODEs, and quantitatively test its solution performance such as accuracy and efficiency.
Table 1: Comparison of the Properties of the Three Methods for Solving Differential Equations.
Method Numerical Discretization Conventional NN PINN
Features:
✓Advantage
$\times$ Disadvantage
✓Complete theoretical
foundation, interpretable,
readily available error
estimates, with convergence
and stability guarantees
✓High efficiency
✓High precision
$\times$ Suffering from curse
of dimensionality
$\times$ Mesh generation is
still complex
$\times$ Inability to seamlessly
integrate noisy data
$\times$ Excessive cost of
solving inverse problem
✓Mesh-free
✓Breaking the curse
of dimensionality,
can cope with
high dimensional and
nonlinear problems
✓Generalizability to
new data samples
$\times$ Need big data and
strong computing power
$\times$ Overfitting
$\times$ Black box issue
✓Advantages of
conventional NN +
✓Small data, less prone
to overfitting, strong
generalization capability
✓Robust to imperfect data
and incomplete models
✓Effective and efficient in
dealing with ill-posed
and inverse problems
✓Highly interpretable
✓Fewer network parameters,
and fast training
✓Flexible cross-platform
1.4 Related Work & Our Contributions
[…]
Related Work on PINN Solving Oscillator ODEs Baty et al. [26] carried out a systematic and comprehensive benchmarking of PINN solving typical linear/nonlinear oscillator ODEs based on PyTorch.  [29] addressed the problems of (i) limited known solution data, and (ii) integration intervals that, if too large, would make it difficult for the PINN to accurately predict the solution, when solving initial value problems (IVPs) for stiff ODEs. Improved strategies such as embedding more physical information, optimizing the training data loss to ensure that it considers all initial conditions more comprehensively, and incremental learning strategy, i.e., gradually increasing the integration intervals and optimizing the parameters of the PINN in each interval, and employing a shifting mesh to minimize the residuals of the DEs, have resulted in the improved PINN having a higher accuracy and a more stable training in solving IVPs for stiff ODEs process. In addition, these improved strategies have been shown to be equally effective in solving boundary value problems (BVPs), such as the solution to the high Reynolds number steady-state convection-diffusion equation.  [30] further compared the performance of soft and hard constraints in solving higher-order Lane-Emden-Fowler type equations. The results show that soft constraints are more flexible and can handle many different physical scenarios and boundary conditions, but may require more training data and iterations to converge to a satisfactory solution. Hard constraints are able to satisfy the constraints of the differential equations more directly, and therefore may achieve higher accuracy in fewer iterations, but may require a more complex process of test solution selection and tuning. The PINODE method [31] constructs reduced-order models (ROMs) based on an autoencoder and aids model training with a physics-informed loss term. Using the collocation point technique, the residuals of the physics laws are added to the loss function as regularization terms, thus optimizing both data fitting and physical consistency during the training process. Recently, [32] proposed several improved PINN methods for solving second-order ODEs with sharp gradients, namely, (i) Deep-TFC via introducing constraint expressions from the theory of functional connectivity (TFC) and satisfying the constraints of the ODE analytically. (ii) PIELM which trains single-layer neural networks using the extreme learning machine (ELM) algorithm. These two methods have higher accuracy and efficiency compared to the vanilla PINN. (iii) X-TFC combines the advantages of Deep-TFC and PIELM to further improve the performance.
[…]
Our Contributions
  • The advantages of PINN over conventional neural networks, traditional numerical discretization methods are surveyed and summarized in solving ODEs and also PDEs through relevant literature iversgitation (see Table 1).
  • We formalize the mathematical framework and computational flow of the soft-constrained PINN approach suitable to solving both ODEs and PDEs. Following and compared to [26], we then explore the working mechanism as well as characteristic performance of the PINN method experimentally, through solving typical oscillator (covering first- and second-order, linear and nonlinear) ODEs. In particular, we focus on the effectiveness, high efficiency, and robustness to noise of the soft-constrained PINN with minimal data. Our DeepXDE-based PINN implementation improves coding and solution efficiency.
  • Our verification comprehensively and systematicly shows the intrinsic working mechanism or logic of PINN that: 1) PINN embeds physical information and prior knowledge to mechanically reduce data redundancy and has the unique advantage of achieving strong generalization with small data. It greatly reduces the need for labeled data and can even predict solutions with minimalistic data (e.g., linear first- and second-order differential equations require only one and two training data containing initial values, plus a few collocation points, respectively), and also improves computational efficiency. 2) We find out that PINN is robust to noisy data and provides accurate and physically consistent predictions, and even extrapolates outside the time domain of the training set, therefore the generalization ability is enhanced. And with 3) better intepretability.
  • Taking the convergence of a second-order nonlinear Duffing oscillator with minimalistic data (2 training points containing initial values plus some collocation points) as a validation example, it is shown that the soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law.
  • The advantages of the DeepXDE-based implementation of the PINN method, such as light code, fast training speed, and flexible across CPU/GPU platforms are demonstrated.
This paper is organized as follows. In 1 Introduction, we point out that the success of deep neural networks relies on big data and strong computational power, and that there are improvements to be made especially in small data scenarios, leading to the PINN method that can leverage prior information and knowledge (1.1). In terms of solving ODEs/PDEs, it is then adequately compared with conventional neural networks and traditional numerical discretisation methods (1.2), thereby distilling the features and advantages of the PINN method (1.3). Then, using solving oscillator ODEs as an analytical and experimental case study, we briefly describe the related work and summarise the contributions of ours (1.4). In 2 Method & Result, 2.1 establishes the mathematical framework and computational flow of the soft-constraint PINN approach focused in this paper and the DeepXDE-based implementation. 2.2 demonstrates the working mechanism of the soft-constraint PINN approach and its accuracy and efficiency in details with examples of typical oscillator ODEs (covering first-/second- order, linear/nonlinear): 1) PINN greatly reduces the need for labeled data, works even with minimalistic data, speeds up training, and can also predict data outside the time domain of the training set (2.2.1). 2) The nonlinear case is consistent with the linear case (just need to add extra training points), and PINN has good robustness to noisy data, reflecting its strong generalization performance. DeepXDE-based implementation of PINN is light code with high training efficiency, and flexible across GPU/CPU platforms; the acceleration effect of GPU is obvious, however, the small data characteristics of PINN make it also very efficient to train on CPU (2.2.2). In particular, the comparison of convergence results of second-order nonlinear Duffing oscillator with or without the regularization of conservation of energy under minimalistic data demonstrates that soft-constrained PINN can easily integrate the physical laws to improve the solution results (2.2.3). Finally, in 3 Concluding Remarks, this work is summarized in 3.1 and 3.2 outlooks some future steps.
2 Method & Result
2.1 Soft-constrained PINN & Implementation
2.1.1 Governing Equations
Consider a physical system defined on a spatial or spatio-temporal domain $\Omega\subseteq\mathbb{R}^{d}$, where the unknown $u(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is system state variable that are functions of spatial or temporal coordinates $\boldsymbol{x}\in\Omega$. For time-independent systems, $\boldsymbol{x}=(x_{1},\ldots,x_{d})$; for time-dependent systems, $\boldsymbol{x}=(x_{1},\ldots,x_{d-1},t)$. The physical laws of the dominant system are often characterized in terms of ODEs/PDEs, and these equations are known as the governing equations [33], the general form of which is given by
[…]
For time-dependent systems (i.e., dynamical systems), initial conditions $\mathcal{I}(u;\theta)(x,t_{0})=0,x\in\Omega_{0}$ (Eq. (2)) need to be set for the state variables (and sometimes their derivatives) at the initial moment $t_{0}$. $\theta\in\Theta$ parameterizes or controls the system, where $\theta$ can be a vector or a function included in the control equation. For systems characterized by PDEs, we also need to constrain the state variables on the boundaries of the spatial domain $\partial\Omega$ in order to make the system well-posed. For the boundary points $x\in\partial\Omega$, we have boundary conditions $\mathcal{B}(u;\theta)(x,t)=0,x\in\partial\Omega$ (Eq. (3)). If there are no constraints on the initial and boundary conditions, then $\mathcal{I}(u;\theta)\triangleq 0$ and $\mathcal{B}(u;\theta)\triangleq 0$ [33].
2.1.2 Framework of PINN via Soft Constrains
As shown in Figure 1, suppose there is a system obeying Eq. (1) and a dataset $\{u(\boldsymbol{x}_{i})\}_{i=1,\ldots N}$. It is then possible to construct the neural network $u_{w}(\boldsymbol{x})$ and train it with the following loss function, namely
$\begin{split}\mathcal{L}_{total}=\mathcal{L}_{data}+\underbrace{\mathcal{L}_{gov}+\mathcal{L}_{IC}+\mathcal{L}_{BC}+\cdots}_{\mathcal{L}_{phys}}\\ =\frac{\lambda_{d}}{N}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{g}}{|\Omega|}\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\\ \frac{\lambda_{i}}{|\Omega_{0}|}\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\frac{\lambda_{b}}{|\partial\Omega|}\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}+\cdots.\end{split}$ (4)
Where, $\mathcal{L}_{data}=\frac{1}{N}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}$ is the regular data loss of PINN in matching the labled datasets; $\mathcal{L}_{gov}=\int_{\Omega}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ is the residual loss that makes the network $u_{w}$ satisfy the governing equation constraints; $\mathcal{L}_{IC}=\int_{\Omega_{0}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ and $\mathcal{L}_{BC}=\int_{\partial\Omega}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x})\|^{2}\mathrm{d}\boldsymbol{x}$ are the corresponding losses that make PINN satisfy the initial and boundary conditions, respectively.
PINN losses are flexible and scalable: The $\cdots$ in Eq. (4) can be additional regularization terms $\mathcal{L}_{reg}$, e.g., those satisfying the law of conservation of energy; If there are no available data or initial/boundary constraints, the corresponding loss terms can simply be omitted. The learning weights of these losses can be set by adjusting the hyperparameters $\lambda_{d}$, $\lambda_{g}$, $\lambda_{i}$, $\lambda_{b}$, $\cdots$.
In order to compute (4), several integral terms need to be evaluated, which involve the computation of higher-order derivatives of $u_{w}(\boldsymbol{x})$. PINN utilizes automatic differentiation of computational maps to calculate these derivative terms. The integrals are then approximated using a set of collocation points, which can be sampled using the Monte-Carlo method. We use $N_{d}$, $N_{g}$, $N_{i}$, $N_{b}$ to denote the number of corresponding data or collocation points. Then, the loss function can be approximated as [33]
$\begin{split}\mathcal{L}_{total}=\frac{\lambda_{d}}{N_{d}}\sum_{i=1}^{N}\|u_{w}(\boldsymbol{x}_{i})-u(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{g}}{N_{g}}\sum_{i=1}^{N_{g}}\|\mathcal{F}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\\ \frac{\lambda_{i}}{N_{i}}\sum_{i=1}^{N_{i}}\|\mathcal{I}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\|\mathcal{B}(u_{w};\theta)(\boldsymbol{x}_{i})\|^{2}+\cdots.\end{split}$ (5)
Eq. (5) can be efficiently trained using first-order optimizers such as SGD and second-order optimizers such as L-BFGS.
Figure 1: Schematic Diagram of PINN via Soft Constrains.
2.1.3 Implementation Based on DeepXDE
We implemented soft-constrained PINN with DeepXDE [25] and compared the results with a PyTorch-based implementation [26] (see  2.2.2 for details). DeepXDE is well-structured and highly configurable. Code written in DeepXDE is shorter, closer to mathematical formulas, and more computationally efficient. Solving differential equations in DeepXDE is akin to “building blocks” using built-in modules that specify the computational domain (geometric and temporal), differential equations, boundary/initial conditions, constraints, training datasets, neural network structure, optimization algorithms, and hyper-parameters, etc., all of which are loosely coupled.
2.2 Results of Soft-constrained PINN Solving Oscillator ODEs
For direct comparison, we choose to solve the same linear and nonlinear oscillator ODEs (i.e., Tutorial Example and Van der Pol osillators) as in [26].
2.2.1 PINN vs NN to Solving Linear Oscillator ODE
The linear oscillator ODE (which we named the Primer oscillator) is
$\frac{du}{dt}+0.1u-\sin(\pi t/2)=0.$ (6)
Where, $t\in[0,30]$, initial condition $u_{0}=1$. We solved it with conventional NN and PINN respectively. Both used a 3-layer 32-neuron ($3\times 32$) fully connected hidden layer as base network, Adam optimizer, learning rate $\eta=3\times 10^{-3}$. The exact solution was given by the result obtained by the Runge-Kutta method (3000 steps, at least about 200 steps).
When solving with NN, if the number of training points is insufficient or not well distributed, for example: 1) Case 1 in Figure  2(a)  samples 26 training points uniformly in the entire time domain and does not converge well; 2) Case 2 in Figure  3(a)  samples 61 training points uniformly in the left half of the time domain. Although the left half of the interval is well fitted, it may fail completely on the right half of the interval without training points, i.e., there is no extrapolation capability. In this case, at least 50 training points need to be sampled uniformly over the entire time domain to get a better match with the exact solution.
The only difference between solving with the soft-constrained PINN approach in this example and a conventional NN is the addition of $\mathcal{L}_{phys}$ to the total loss function. Since the initial condition in the conventional NN is already included in the training set (i.e., the first training point on the left), only $\mathcal{L}_{gov}$ satisfying Eq. (6) is added to total loss. This is accomplished by adding extra collocation points (see Eq. (5) ), in this example sampling at least 50 collocation points uniformly throughout the entire time domain. It could be understood that the training points are supervised and the collocation points are unsupervised. The exact number of collocation points required depends on the number of neural network parameters, i.e. the size of the problem being solved. The distribution of collocation points (uniform/random) also has an impact on the results [27]. The learning rate and loss weights also affect the convergence of the gradient descent algorithm [26]. Here the loss weights were taken as $\lambda_{d}=1.0,\lambda_{r}=6\times 10^{-2}$ through hyperparameter optimization. The loss weights act as a balance between the dual drivers of data and physical information.
For the two cases where NN fails to converge, the results obtained by the PINN method are shown in Fig. 2(b) and Fig. 3(b), respectively, and a significant improvement can be seen when comparing with the corresponding figure (a) on the left: 1) The number of supervised training data points required is reduced from 50 to 26 (the minimalist case in this example actually requires only the initial value point plus another 48 unsupervised collocation points, see Fig. 4). In contrast, the Runge-Kutta method needs at least about 200 time steps due to the stability and accuracy requirements. Thus the need for both supervised and unsupervised sampling points is drastically (even orders of magnitude) reduced by the PINN method - one of the significant advantages of PINN. 2) In particular, the generalization ability is enhanced by the assistance of collocation points and the use of physical information carried by ODEs, which equips PINN with the ability to extrapolate data outside the time domain of the training set. In addition, once PINN has completed its training, the prediction for new data sample is instantaneous, which is a property that numerical discretization (e.g., Runge-Kutta) methods do not have.
(a) $N_{data}=26$, $N_{c}=0$
(b) $N_{data}=26$, $N_{c}=50$
Figure 2: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 1. $N_{data}$–No. of training data, $N_{c}$–No. of collocation points, all the same below. (Reproduced from [26] via DeepXDE for comparison.)
(a) $N_{data}=60$ (ONLY Left Interval), $N_{c}=0$
(b) $N_{data}=60$ (ONLY Left interval), $N_{c}=50$
Figure 3: Result Comparison of PINN vs NN Solving Linear Oscillator–Case 2. (Reproduced from [26] via DeepXDE.)
[…]
Minimalistic Training Data So, what is the minimum amount of training data required for the PINN method? For this example, if only one point, the initial value, is taken and the collocation points are kept to be 50, a good solution is obtained, as shown in Fig. 4(a), and the loss function descent process is shown in Fig. 4(b). [26] has already showed that when the nonlinearity of the problem is weak, a very small amount of labeled training data plus a few collocation points is sufficient to predict the solution. In the minimalist case, just as classical analytic or numerical integration methods require only one and two initial conditions for solving first- and second-order differential equations, respectively. PINN also require only one training point (initial value) and two training points containing the initial value point, respectively (see Fig. 7(b)).
(a) $N_{data}=1$, $N_{c}=48$
(b) Corresponding history of the loss function
Figure 4: Minimalistic Training Data Example of First-order Linear ODE. (Reproduced from [26] via DeepXDE.)
[…]
Accelerating Training Time Compared to NN, PINN can significantly shorten the training time because it requires only a small amount of data (minimal labeled training data plus a small number of collocation points). On the other hand, as its loss function becomes more complex with the addition of $\mathcal{L}_{phys}$, it requires more computations and sometimes even more iteration epoches when training, which in turn drags down the training efficiency. We performed a validation test of Primer oscillator on DeepXDE 1.11.1 based on a CPU platform (Windows10, Intel Core i9-9900K @3.6GHz, 32GB Memory). The training time is the average of 5 training sessions (24,000 epoches per session, usually 24,000 epoches are not needed before convergence, and 24,000 epoches are trained for fair comparison), and the results are shown in Fig. 5. It can be seen that, for linear oscillator ODEs, the training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. It should be emphasized that since Primer oscillator is a simple linear ODE, we used the same base network ($3\times 32$) for ease of comparison. PINN can accelerate the training even further by reducing the number of network parameters, because PINN need less network parameters than NN due to its small data characteristics. For example, for Fig. 5(b), if the base network is reduced to $3\times 16$, it also converges well while the training time is reduced to 15.30s.
(a) $N_{data}=50$, $N_{c}=0$
(b) $N_{data}=1$, $N_{c}=50$
Figure 5: Training Time of PINN vs NN Solving Primer Oscillator via DeepXDE: (a) 16.56s by NN with Min. 50 training points; (b) 15.23s by PINN with Min. 1 initial value training point + 50 collocation points. (Reproduced from [26] via DeepXDE.)
2.2.2 PINN Solving Non-linear Oscillator ODE
The soft-constrained PINN approach also performs well in solving nonlinear oscillator ODEs [26]. We successfully solved the typical strongly nonlinear Van der Pol (VDP) oscillator based on DeepXDE, and further demonstrated the tolerance of PINN to noisy data as well as the light-code and training speed advantages of the DeepXDE implementation of PINN. The ODE of VDP is
$\frac{d^{2}u}{dt^{2}}+\omega_{0}^{2}u-\epsilon\,\omega_{0}(1-u^{2})\frac{du}{dt}=0.$ (7)
Where, $t\in[0,1.5]$, initial condition $u_{0}=1$; $\omega_{0}$ is the normalized angular velocity, taken as $\omega_{0}=15$; $\epsilon$ reflects the degree of nonlinearity of the VDP, and the larger the value of $\epsilon$ the stronger the nonlinearity.
[…]
Noisy Data Tolerated The PINN method is used to solve the nonlinear VDP oscillator with small and noisy data. As with solving the Primer oscillator, the fully connected base network of $3\times 32$ was still used, which reflects the strong expressive power of neural networks inherited by PINN. The relevant parameter settings (i.e., $\epsilon$, number of training data points, number of collocation points, normal noise variance) are illustrated in Fig. 6, and other hyperparameter settings can be found in the code. Exact solution was generated by the 4th-order Runge-Kutta method. As can be seen from Fig. 6: 1) With the same amount of training data and collocation points, the results obtained from training with noisy data (right column) are basically identical to the noiseless results (left column), showing the good inclusiveness of the PINN method against noise. 2) As the degree of nonlinearity increases (by increasing $\epsilon$), more training and collocation points are needed to be arranged in regions with significant nonlinear features to obtain an exact solution. When $\epsilon=1,3,5$, the least number of training and collocation points are $N_{data}=28,32,38;N_{c}=15,25,40$, respectively. And 3) as the nonlinearity increases, the tolerance to noise decreases: when $\epsilon$ increases from 1 to 5, the $\sigma$ that can be tolerated decreases from 0.1 to 0.05; even increasing the number of data points has a limited effect on improving noise tolerance.
(a) $\sigma=0;N_{data}=28$, $N_{c}=15$
(b) $\sigma=0.1;N_{data}=28$, $N_{c}=15$
(c) $\sigma=0;N_{data}=32$, $N_{c}=25$
(d) $\sigma=0.08;N_{data}=32$, $N_{c}=25$
(e) $\sigma=0;N_{data}=38$, $N_{c}=40$
(f) $\sigma=0.05;N_{data}=38$, $N_{c}=40$
Figure 6: Results of PINN Solving VDP Oscillator via DeepXDE with Noisy and Small Data.
[…]
PINN Implementation Code via DeepXDE vs PyTorch Comparison of Code  1 and Code  2 shows that the overall process (seen in Fig. 1) of realizing soft-constrained PINN based on DeepXDE or PyTorch is consistent. DeepXDE has a better customized encapsulation of the PINN method and thus requires less coding. With the same setup for solving VDP, the number of code lines (about 40 lines) via DeepXDE implementation is roughly one-third as the PyTorch implementation. In addition, the training efficiency of DeepXDE is significantly better than that of PyTorch, thanks to the optimization of DeepXDE for PINN [25]. For the example in Figure 5(b), PyTorch-based training took 43.10s with the same settings, which is about 1.8 times more than that of DeepXDE.
1 import torch
2 import numpy as np
3 from scipy.integrate import odeint
4
5 def VDP:
6 """Define the ODE of VDP"""
7
8 # Define time domain and boundary conditions
9 t = np.linspace(0, 1.5, 1000)
10 y0 = [1, 0] # y0 is the initial condition
11
12 # Call odeint in Scipy to generate numerical exact solutions
13 y = odeint(VDP, y0, t)
14
15 # Sample and generate the training set from the exact solution, including initial point (t0, y0) and endpoint
16 t_data=t[0:1.5:0.005]
17 y_data=y[0:1.5:0.005]
18
19 # Select collocation point
20 x_collocation = torch.linspace(0.,1.5,48).view(-1,1).requires_grad_(True)
21
22 class FNN:
23 """Define neural networks"""
24
25 model = FNN(1, 1, 32, 3) # Instantiate FNN
26
27 # The training process of PINN:
28 for i in range(24000):
29
30 # Data loss
31 yh = model(t_data)
32 Loss_data = 1 * torch.mean((yh - y_data)**2)
33
34 # Physical loss
35 yhp = model(x_collocation)
36 dx = torch.autograd.grad(yhp, x_collocation, torch.ones_like(yhp),
37 create_graph=True)[0]
38 dx2 = torch.autograd.grad(dx, x_collocation, torch.ones_like(dx),
39 create_graph=True)[0]
40 physics = (dx2 + 15 **2 * yhp - 5 * (1 - yhp**2) * dx)
41 Loss_phys = (1e-4) * torch.mean(physics**2) * 1
42
43 # Calculate the mean square error of the test set
44 yhpp = model(t)
45 mse = torch.mean((yhpp - y)**2)
46
47 # Total loss of backpropagation
48 Loss_total = Loss_data + Loss_phys
49 Loss_total.backward()
Listing 1: PyTorch Implementation of PINN to Solve VDP
1 import deepxde as dde
2 import numpy as np
3 from scipy.integrate import odeint
4
5 def VDP:
6 """Define the ODE of VDP"""
7
8 # Call odeint in Scipy to generate numerical exact solutions
9 y = odeint(VDP)
10
11 # Sample and generate the training set from the exact solution, including initial point (t0, y0) and endpoint
12 t_data=t[0:1.5:0.005]
13 y_data=y[0:1.5:0.005]
14
15 geom = dde.geometry.TimeDomain(0.0, 1.5) # Define time domain
16
17 def boundary(t, on_boundary):
18 return on_boundary and dde.utils.isclose(t[0], 0)
19
20 def error_derivative(inputs, outputs, X):
21 return dde.grad.jacobian(outputs, inputs, i=0, j=None)
22
23 # Define initial conditions: y(0)=1 and y’(0)=0
24 ic1 = dde.icbc.IC(geom, lambda x : 1, lambda _, on_initial: on_initial)
25 ic2 = dde.icbc.OperatorBC(geom, error_derivative, boundary)
26
27 # Define training set and boundary conditions (including boundary conditions)
28 t_y_data = dde.PointSetBC(t_data, y_data)
29
30 data = dde.data.TimePDE(geom, ode, [ic1, ic2, t_y_data], num_domain=40, num_boundary=1) # num_domain is the number of collocation points
31
32 net = dde.maps.FNN([1] + [32] * 3 + [1]) # Instantiate FNN
33 model = dde.Model(data, net)
34 model.train(epochs=24000)
Listing 2: DeepXDE Implementation of PINN to Solve VDP
[…]
PINN Computational Efficiency @CPU vs @GPU The PINN method does not depend on big data so it can be efficiently computed on platforms such as CPU or GPU. We compared the training time for solving Primer, VDP oscillators based on DeepXDE on CPU and GPU (CPU as before, GPU is Nvidia GTX1080Ti) respectively, taking the average of 5 training sessions (24000 epoches each session). Base networks use $2\times 32$ and $3\times 32$ fully connected hidden layers, respectively. As shown in Table  2, the training acceleration effect of GPU on PINN solving Primer and VDP is obvious: the training duration is reduced from 16.02s and 17.60s to 8.73s and 13.15s, respectively. When the base network is reduced to $3\times 16$, the training duration results show a consistent trend.
Table 2: PINN Training Time (Second) @CPU vs @GPU.
PINN to ODEs via DeepXDE @CPU @GPU
Primer Oscillator ($N_{data}=50,N_{c}=48$) 16.02 8.73
VDP Oscillator ($N_{data}=50,N_{c}=48$) 17.60 13.15
[…]
2.2.3 Conservation of Energy Regularization Improves Result of Solving Duffing Oscillator
A major advantage of PINN is that it can fully utilize physical information or knowledge as a priori (e.g., the law of conservation of energy). The soft-constrainted approach is implemented by adding the regularization term $\mathcal{L}_{reg}$ to the total loss function. We verified this with a second-order nonlinear Duffing oscillator. The ODE is
$\frac{d^{2}u}{dt^{2}}+\alpha u+\beta u^{3}=0.$ (8)
Where, $t\in[0,1.5]$, take $\alpha=1.0,\beta=1.0$ with initial condition $u_{0}=15,u^{\prime}_{0}=0$. Again, the exact solution was obtained by Runge-Kutta method. The conservation of energy term of Eq. (8) is $E=\frac{1}{2}(\frac{du}{dt})^{2}+\frac{1}{2}\alpha u^{2}+\frac{1}{4}\beta u^{4}$, substituting to the regularization term $\mathcal{L}_{reg}$.
Considering the minimalistic data case, for second-order ODEs only two training points containing initial values were used, and the number of collocation points $N_{c}=40$ was chosen uniformly. The optimal tuning results of the PINN solving Duffing oscillator with and without the conservation of energy regularization term are shown in Fig. 7(a) and (b), respectively. The PINN without the conservation of energy regularization in the minimalistic data case, which is affected by nonlinearity, fails to complete convergence throughout the entire 72000 epoches of training; In contrast, the PINN with the conservation of energy regularization achieves a significant improvement, which is in perfect agreement with the exact solution. It should be noted that the conservation of energy regularization is useful for Duffing because Duffing follows the conservation of energy; It is useless for VDP because VDP is energy dissipative.
(a) w/o the Conservation of Energy Regularization
(b) with the Conservation of Energy Regularization
Figure 7: Results of PINN Solving Duffing Oscillator with the Conservation of Energy Regularization or Not.
3 Concluding Remarks
3.1 Summary
(1) The success of deep neural networks relies on big data and strong computational power, yet neglects to fully utilize prior information or knowledge. In particular, it lacks robustness or even fails to converge in the case of small data. PINN is precisely a powerful complement to deep neural networks in small data scenarios. It inherits the strong expressive ability of neural networks and incorporates physical information and knowledge into neural network’s topology or learning architecture as well as computational process by means of hard-coding or soft constraints, which reduces data redundancy, improves generalization performance, and consequently increases computational efficiency, robustness, and interpretability. More specifically, PINN generalizes well even with only a small amount of (supervised or labeled) training data, remains robust to data imperfections (e.g., noise, missing or outliers), and provides accurate, fast and physically consistent predictions that can even be used for extrapolation tasks.
On the other hand, compared to traditional numerical discretization methods (e.g., Runge-Kutta method, FEM), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods. Thus PINN is becoming a favorable catalyst for the emerging era of Digital Twins [34, 35, 28, 36, 37].
(2) We generalized the mathematical framework and computational flow of soft-constrained PINN for solving ODEs/PDEs. It is clear from the composition of the loss function that PINN is driven by both data and physical information. Compared to conventional NN, PINN adds differential equations (including initial, boundary conditions) and a priori information or knowledge (e.g., the law of conservation of energy) as constrains or regularization. Following and compared to [26], the working mechanism of the soft-constrained PINN and its accuracy and efficiency were experimentally verified by solving typical (covering first- and second-order, linear and nonlinear) oscillator ODEs such as Primer, Van der Pol, and Duffing oscillators.
PINN greatly reduces the need for labeled data. When the nonlinearity of the oscillator ODEs (e.g., Primer, Duffing) is weak, a very small amount of supervised training data plus a few unsupervised collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (including initial values) are needed for first- or second-order ODEs, respectively. This is just as classical analytic or numerical integration methods require only one and two initial conditions, respectively, for solving first- and second-order differential equations. Strongly nonlinear oscillator ODEs (e.g., VDP) also require only an appropriate increase in the number of training points, which still has significant advantages over conventional NN. With the aid of collocation points and the use of physical information or knowledge, PINN has the ability to extrapolate data outside the time domain of the training set, and is robust to noisy data, thus with enhanced generalization capabilities. The training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. Small data allows fewer network parameters, so it can be further accelerated.
The second-order nonlinear Duffing oscillator regularized by conservation of energy is able to converge quickly with minimalist data (2 training points containing initial values plus a few collocation points), whereas it does not converge without the conservation of energy regularization. The case study results show that, soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law.
(3) The DeepXDE-based implementation of PINN is not only light code but also efficient in training. The acceleration effect of GPU on PINN is obvious, as PINN does not depend on big data, it has high cross-platform flexibility and can also be efficiently calculated on CPU platform.
3.2 Future Work
Based on this work, the next step could be to continue the exploration of PINN methods (both hard-constrained and soft-constrained) for solving stiff ODEs and PDEs, as well as the exploration of sampling methods [27, 28], which are critical to the accuracy and efficiency of the solution. In addition, PINN combined with domain decomposition can be extended to large problems [8, 38], and further extended to solve integral differential equations (IDEs), fractional order differential equations (FDEs) and stochastic differential equations (SDEs) [25].
[…]
Acknowledgements The authors would like to thank Prof. Hubert Baty from University of Strasbourg for his kind instruction and discussion. We would also like to thank the anonymous reviewers for their comments and suggestions.
[…]
Funding This work was sponsored by the National Natural Science Foundation of China (grant No. 51405289), and funded by Research Programm of Shanghai Jian Qiao University (KYJF21XX202122).
[…]
Conflict of interest The authors declare no conflict of interest.
[…]
Data/Code availability The data in the paper are computationally generated from the experimental codes. The codes will be available at https://github.com/mikelu-shanghai/PINNtoODEwithSmallData. (Readers in need can contact us.)
[…]
Author contribution Conceptualization & funding acquisition, Lu K.L. and Bi Z.; methodology & software & writing—original draft preparation, Lu K.L. and Su Y.M.; resources & investigation & validation, Lu K.L., Bi Z., Su Y.M. and Qiu C.; writing—review and editing, Lu K.L., Su Y.M. and Zhang W.J.; supervision, Zhang W.J.; All authors have read and agreed to the published version of the manuscript.
Mathematical Notations
Notations Description
$u$ state variables of the physical system
$\boldsymbol{x}$ spatial or spatial-temporal coordinates
$x$ spatial coordinates
$t$ temporal coordinates
$\theta$ parameters for a physical system
$w$ weights of neural networks
$\mathcal{F}$ differential operator representing the PDEs/ODEs/etc.
$\mathcal{I}$ initial conditions (operator)
$\mathcal{B}$ boundary conditions (operator)
$\Omega$ spatial or spatial-temporal domain of the system
$\Theta$ space of the parameters $\theta$
$W$ space of weights of neural networks
$\mathcal{L}_{total}$ total loss function
$\mathcal{L}_{data}$ supervised data loss
$\mathcal{L}_{gov}$ governing equation residual loss
$\mathcal{L}_{BC}$ boundary condition loss
$\mathcal{L}_{IC}$ initial condition loss
$\mathcal{L}_{reg}$ regularization loss
[…]